# Autonomation, not Automation: Activities and Needs of Fact-checkers as a Basis for Designing Human-Centered AI Systems

### v2

@@ 2. Related work
The lack of end-user viewpoint in the early design-related activities is well known from the Human-centered Design (HCD) practice (Hartikainen et al., 2022). Specifically, while tech-savvy fact-checkers are often called for testing of developed AI tools, these often turn out to be “irrelevant” for their work (Dierickx, 2022). This can be prevented by early inclusion of fact checkers in the tool design process. The challenge is that the capabilities of AI are unclear to users who set the end-user requirements  (Hartikainen et al., 2022). This disproportion calls for an interdisciplinary approach, where the practices and problems of users (fact-checkers) are studied by social science and humanities (SSH) researchers and at the same time discussed with AI researchers to design powerful AI tools.
@@ 3. Methodology
Human-centered artificial intelligence (HCAI) is based on the processes that extend user experience design methods such as stakeholder engagement (Shneiderman, 2022). The goal is to create tools that augment and enhance human performance. HCAI systems emphasize a high level of human control, while embedding high levels of automation that can be achieved by a good design (Shneiderman, 2022).
Human-Centered Design (HCD) approaches are mentioned to be capable of contributing to the field of HCAI as well (Auernhammer, 2020). Specifically, Design Thinking, previously named Need-Design Response (NDR) focuses on the design and development of any tools or systems for the physical, intellectual, and emotional needs of people. It allows identifying human needs in the early phases of the design project through practices such as need finding.
Human-Centered Interaction (HCI), as an interdisciplinary field, adopts a ’human-centered design’ approach to develop computing products that meet user needs, which makes this field a potentially strong contributor to HCAI. Nevertheless, the development of AI systems is still mainly driven by a “technology-centered design” approach  (Shneiderman, 2022; Xu, 2019). To respond to the challenge, a human-centered AI approach was proposed that places humans at the center of AI design as the ultimate decision makers  (Shneiderman, 2022; Xu, 2019).
@@ 3.1. Research design
To analyze the activities and needs of fact-checkers, we first engaged these stakeholders individually semi-structured in-depth interviews. Using the iterative content analysis, we identified the activities — repeating routines performed by fact-checkers in the individual stages of the fact-checking process. These activities were further validated with previous research work on fact-checker practices and visualized by conceptual maps of the fact-checking process.
By proceeding from the content analysis, we also identified particular fact-checkers’ needs. Nevertheless, the needs are to some extent implicit, since they are inner motivational states to reach goals and it is possible to be unaware of one’ s true needs  (Case and Given, 2016). As not all respondents were tech-savvy, they referred more often to problems instead of directly formulating their needs for specific technological support. Therefore, in this paper, we use problems of fact-checkers as a substitution of their needs (i.e., a corresponding need refers to finding a solution that can solve the problem expressed by a fact-checker). The most significant problems (and resulting needs), connected to the activities they relate to, were consequently verified by a quantitative validation survey. Finally, the implications for AI tasks and tools were derived. We illustrate this methodology (research process, methods, and terminology) in Fig. 2.
@@
  • RQ1: Which information resources, procedures, and information technologies do fact-checkers use and how do they depend on local specifics?
@@
To answer RQ1, we conducted semi-structured in-depth interviews, in which we were interested how Central European fact-checkers: 1) monitor the online space to identify potential misinformation; 2) select potential false claims/narratives; 3) communicate and avoid potential duplication; 4) verify content credibility and veracity; and finally, 5) disseminate fact-checks. We compared our findings with fact-checkers worldwide.
@@ 3.2. Selection criteria for respondents
The in-depth interviews were conducted with nine fact-checkers from five major Central European (Slovak, Czech, and Polish) fact-checking organizations. All of the respondents performed fact-checking professionally as part of their full-time job, none of them was employed in social media. We involved both external and in-house professionals in management positions (editors in chief, project coordinator), fact-checking only positions, but also fluid and overlapping fact-checking roles with a role of a journalist, PR manager, editorial manager and senior research fellow. Our sample represents one of the low-resource language groups, under-explored in AI-research. The sample covers a variety of sizes and types of organizations – from small NGOs to large news agencies. The majority of organizations also collaborated with social media providers (e.g., Meta). Most of them were part of the IFCN network, which prohibits any kind of political connection. Two organizations were not members of the IFCN (and did not collaborate with social media). These organizations focused on the fact-checking of political discussions in mass media (mainly TV) and partisan news. The product of these two organizations was an article, summarizing the main arguments in a misleading topic of interest (not a structured fact-check of a claim).
For this work, we define Central Europe as the countries of the Visegrád Group (Tiersky et al., 2004). While Central Europe geographically belongs to the European region, its historical and linguistic context differs from Western Europe (eastern block legacy, Slavic languages). This has many consequences on its current social and cultural setting and related issues (e.g., the level of Russian influence). Contextual factors affect the choices of action and the use of sources and channels in the online environment (Agarwal, 2018). We can thus assume that some needs of Central European fact-checkers can differ from the fact-checkers elsewhere in Europe/world. In particular, we can expect differences when compared with high-resource language areas.
@@
All fact-checkers participating in the survey had to be members of recognized professional groups. Therefore, organizations from CEDMO (Central European Digital Media Observatory) hub were selected for the in-depth interviews. The survey was directed at EDMO (European Digital Media Observatory) hubs members and to the International Fact-checking Network (IFCN) signatories operating in Europe. The survey respondents were recruited through contacts provided by the EDMO or LinkedIn Premium account. Both external fact-checkers and in-house professionals were involved, but none of them was employed in social media. The position names of the respondents mentioned in their LinkedIn profiles were fact-checker, verification officer, journalist, redactor, or analyst in the fact-checking parts of the organizations.
This research involved human participants. Research planning, conduct and reporting was consistent with local regulatory laws and regulations and aligned with ethical principles, such as the ACM Code of Ethics and Professional Conduct and international and national standards for such research.
@@ 3.3. Method 1: Interviews with fact-checkers
The semi-structured interviews were conducted with one organization at a time and each required between one and one and a half hours. The interviews were conducted from December 2021 to February 2022 using a video conference system using the English, Slovak, or Czech language.
When designing the interview, we faced many ambiguities in the fact-checking process. The number of stages, as well as terms used to name them, varies across the literature and tends to depend on the view that the research is dedicated to (social sciences versus AI) as shown in Fig. 3. Existing work in AI typically recognizes four core fact-checking stages, while the first study of today’s practices of fact-checkers (noa, 2020) mentions three different stages. Therefore, we were obliged to redefine the stages found in the literature when designing the interview questions.
@@ 3.4. Method 2: Validation survey
The interviews yielded many problems as well as some implicit inputs on what is desired and needed for fact-checkers to be more efficient in tackling disinformation. These inputs were discussed, using brainstorming methods, and problems that were most cognitively demanding, tedious, time demanding, and complex, as well as brought inconveniences and uncertainty with them were selected. Finally, 13 problems were chosen for the validation survey based on the user activities and current technical feasibility, resource feasibility, and sustainability of potential AI solutions (tools) (see the validation survey questions in Annex B).
The survey also served to validate the individual responses, therefore for each activity, the respondents were queried on:
  • how often fact-checkers need to be involved in such activity (on a six points Likert scale ranging from “more times a day”=6 to “never”=1) and
  • how difficult the activity is (five points Likert scale ranging from “No problem, I like to do it”=1 to “I perceive it as a big problem. I really need help with this”=5).
@@
where $C(a_{i})$ is an importance coefficient that quantifyes the need for automation within activity $i$, $med(f_{i})$ is a median score of the frequency of the fact-checkers’ involvement in activity $i$ and $med(l_{i})$ is a median score of the level of difficulty of activity $i$. As such, the most severe problem would get a coefficient of 30, and the least important problem a coefficient of 1.
@@ 3.5. Results presentation
The results of the interviews, i.e., the details of the fact-checking process, based on the fact-checkers’ answers in the interviews, are summarized and visualized with conceptual models using CmapTools. They are also compared with the fact-checking process of the organizations in (Micallef et al., 2022), fact-checkers’ needs in (noa, 2020; Juneja and Mitra, 2022), and utilized tools in (Dierickx, 2022). Differences are outlined in different fonts according to the legend under each figure.
The identified problems may vary across the different tasks that fact-checkers are involved in. This may be considered a limitation of our study and prompts future research focused on complex process analyzes of individual fact-checking organizations.
@@ 4. Results and findings: Interviews with fact-checkers
In general, regarding RQ1, the results of the interviews indicate that the fact-checking process in Central Europe is similar to that in the rest of the world. However, since the fact-checking process differs in the various research studies (see Fig. 3) and some of the stages were missing according the results of our content analysis, we renamed and extended its stages from the previous works (see Fig. 3). Stage 2 (Selection of potential false claims/narratives) and Stage 4 (Verification of content credibility and veracity, in computer science labeled as evidence retrieval and claim verification) were inspired by (Barrón-Cedeño et al., 2020; Nakov et al., 2021a), but based on the interviews with fact-checkers, the names of the stages were renamed (extended) following these findings: (1) fact-checkers do not select only claims, but also narratives, (2) fact-checkers verify the sources of claims besides individual claims (collectively referred to as content).
We expanded the fact-checking process as it has been understood in previous AI research in: Stage 1 (Monitoring the online space), Stage 3 (Communication and avoiding duplication) and Stage 5 (Dissemination of fact-checks). We also extended the dissemination part of the process: no longer it means only the publication and marketing of fact-checks, but includes a novel subcategory - the identification of the other versions (i.e., other appearances in the online content) of the same misinformation. We use the term “misinformation”, as not every respondent was fact-checking at the level of claims: some of them were interested in repetitive topics (narratives), whereas their final product was a fact-checking article (not a structured fact-check of a claim). This is in accordance with two identified types of fact-checking (Juneja and Mitra, 2022): (1) short-term claims centric, and (2) long-term advocacy centric.
As we found out, some identified problems were specific to the investigated region and connected mainly to the capacity and financial reasons (e.g. weaker collaboration with other fact-checkers or dissemination in an attractive form) and limited language capabilities of the tools (e.g. for monitoring potential misinformation or for finding the previously published fact-checks). The virtue of Central European fact-checkers was ingenuity when forced to use the freely available tools. Its manifestation could be seen in the use of personalization on social networks using fake profiles to monitor misinformation, but also in the creative use of CrowdTangle to monitor notorious liars. Below we present a detailed description of the process and problems of our respondents in comparison with the fact-checkers around the world.
@@ 4.1. Monitoring the online space
“Monitoring of online space is done in two ways: manually or automated. When authors browse media manually (the classic media such as New York Times, Washington Post), they search for topics that are interesting for our audience. Some of us also do automated monitoring in CrowdTangle… I do that every morning by coffee, where I check notorious sources - some lists have about 800 [social media] pages.“
@@
Monitoring is a part of the process that fact-checkers are the most open to autonomating (Dierickx, 2022). Still, there is just a limited pool of tools that are a bit helpful for Central European as well as other non-English speaking fact-checkers on social media, such as CrowdTangle, Meta’s fact-checking tool, and Tweetdeck. These tools have many limitations; e.g. CrowdTangle was originally developed for marketing purposes. Its overemphasis on virality is not always useful for fact-checkers. This was mentioned by our respondents as well as other fact-checkers around the world (noa, 2020; Dierickx, 2022; Micallef et al., 2022). Despite this, fact-checkers agreed on the usefulness of this tool for creating and monitoring lists of notorious liars. It is also useful for checking the popularity of misinformation, for example in cases when it was reported by regular users. Fact-checkers also reported their uncertainty about the future of the service, as Meta abandoned its active development and in August 2024 it is about to be deprecated completely without providing a sufficient alternative. TweetDeck (currently branded as X Pro) became in July 2023 a paid proprietary tool that is available only to verified and premium users of the X platform, what reduced its availability and usefulness for fact-checkers. In Norway, fact-checkers rather rely on their Storyboard tool to find problematic content (Dierickx, 2022), in other research  (Juneja and Mitra, 2022) BuzzSumo, Social searcher or Influencer are mentioned.
@@ 4.2. Selection of potential false claims/narratives
Some Central European fact-checkers also mentioned timeliness, factuality, representativeness, controversiality of potential misinformation, or sympathy of the topic to the authors as factors for selection. All these factors are consistent with previous research (Micallef et al., 2022), as well as putting more emphasis on societal impact than virality, including issues that affect minorities or disadvantaged groups. In comparison to our respondents, the volume of reader requests, verifiability, and the importance of providing a different perspective for their readers were mentioned as important factors for prioritization (Micallef et al., 2022). Some respondents in this research preferred to fact-check influencers, while others were aware of the high probability of a backlash from their followers and preferred not to check. Note that fact-checkers all over the world face legal actions. Some even reported facing personal danger, even physical attacks (noa, 2020; Juneja and Mitra, 2022). The latter has not been the case of our respondents, but it underlines a need for a tool that alerts fact-checkers before something goes viral.
Checking whether a claim is verifiable is one of the tasks that the fact-checkers are open to automating the most (Dierickx, 2022). In some countries, it is performed by a simple Google search or by using freely available tools like ClaimBuster (Hassan et al., 2017), BSDetector (Micallef et al., 2022) or Full Fact tool  (Dierickx, 2022). There are also open-source tools developed and used by fact-checking organizations in Latin America that: 1) differentiate between the opinion and factual statements (Beers, 2019); and 2) optimize the process of collecting and responding to WhatsApp reader requests (Bhuiyan et al., 2020). None of these tools was mentioned or utilized by our interviewees. (Dierickx, 2022) mentions five reasons for not adopting AI tools by fact-checkers: 1) lack of knowledge about the existence of such tools; 2) lack of time to learn the tool; 3) lack of resources to pay for the tool; 4) lack of knowledge of how the AI tool works; and 5) difficulty in integrating the tool into the organization’s workflow. Keeping fact-checkers in the design process of the AI tools and appropriate training may help overcome these barriers (Dierickx, 2022).

### v3

@@ 2.2. AI support of fact-checkers
This problem is not common only to fact-checkers. Although the conceptual foundations of HCAI are extensively discussed in recent literature (Shneiderman, 2022; Xu, 2019) and guidelines for building human-centered AI products exist (e.g. (Google, 2021)), the industry practices and methods appear to lag behind (Hartikainen et al., 2022; Bingley et al., 2023). The lack of end-user viewpoint in the early design-related activities is well known from the Human-centered Design (HCD) practice (Hartikainen et al., 2022). Specifically, while tech-savvy fact-checkers are often called for testing of developed AI tools, these often turn out to be “irrelevant” for their work (Dierickx, 2022). This can be prevented by early inclusion of fact-checkers in the tool design process. The challenge is that the capabilities of AI are unclear to users who set the end-user requirements (Hartikainen et al., 2022). This disproportion calls for an interdisciplinary approach, where the practices and problems of users (fact-checkers) are studied by social science and humanities (SSH) researchers and at the same time discussed with AI researchers to design powerful AI tools.
@@ 3. Methodology
In this study, we aim to address the identified gap in the field – our aim is to investigate the activities and needs of fact-checkers to support research and development of the effective and appropriate human-centered AI tools to counter false information. We specifically focus on the European context, which (as shown in Section 2.1) has multiple historical, cultural, language, and legislative specifics; and remains under-studied in this area. Furthermore, the analysis of existing works as well as practical deployment of AI-based tools for fact-checkers showed a need for more human-centered AI approaches.
Human-centered AI (HCAI) is based on processes that extend user experience design methods such as stakeholder engagement (Shneiderman, 2022). The goal is to create tools that augment and enhance human performance. HCAI systems emphasize a high level of human control while embedding high levels of automation that can be achieved by a good design (Shneiderman, 2022).
Human-Centered Design (HCD) approaches are mentioned to be capable of contributing to the field of HCAI as well (Auernhammer, 2020). Specifically, design thinking, previously named Need-Design Response (NDR), focuses on the design and development of any tools or systems for the physical, intellectual, and emotional needs of people. It allows identifying human needs in the early phases of the design project through practices such as need finding. User needs, especially information needs and user interaction with information, focused primarily on the cognitive viewpoint are studied also in information science, specifically in information behavior studies (Bates, 2009).
Human-Centered Interaction (HCI), as an interdisciplinary field, adopts a “human-centered design” approach to develop computing products that meet user needs, which makes this field a potentially strong contributor to HCAI. Nevertheless, the development of AI systems is still mainly driven by a “technology-centered design” approach  (Shneiderman, 2022; Xu, 2019). To respond to the challenge, a human-centered AI approach was proposed that places humans at the center of AI design as the ultimate decision makers  (Shneiderman, 2022; Xu, 2019).
@@ 3.1. Research design
To analyze the activities and needs of fact-checkers, we first engaged these stakeholders individually in semi-structured in-depth interviews. Using the iterative content analysis, we identified the activities – repeating routines performed by fact-checkers in the individual stages of the fact-checking process. These activities were further validated with previous research work on fact-checker practices and visualized by conceptual maps of the fact-checking process.
By proceeding from the content analysis, we also identified particular fact-checkers’ needs. Nevertheless, the needs are to some extent implicit, since they are inner motivational states to reach goals, and it is possible to be unaware of one’ s true needs  (Case and Given, 2016). As not all respondents were tech-savvy, they referred more often to problems instead of directly formulating their needs for specific technological support. Therefore, in this paper, we use problems of fact-checkers as a substitution for their needs (i.e., a corresponding need refers to finding a solution that can solve the problem expressed by a fact-checker). The most significant problems (and resulting needs), connected to the activities they relate to, were consequently verified by a quantitative validation survey. Finally, the implications for AI tasks and tools were derived. We illustrate this methodology (research process, methods, and terminology) in Fig. 2.
@@
  • RQ1: Which information resources, procedures, and information technologies do fact-checkers use, and how do they depend on local specifics?
@@
To answer RQ1, we conducted semi-structured in-depth interviews, in which we were interested in how Central European fact-checkers: 1) monitor the online space to identify potential misinformation; 2) select potential false claims/narratives; 3) communicate and avoid potential duplication; 4) verify content credibility and veracity; and finally, 5) disseminate fact-checks. We compared our findings with fact-checkers worldwide.
@@ 3.2. Selection criteria for respondents
The in-depth interviews were conducted with nine fact-checkers from five major Central European (Slovak, Czech, and Polish) fact-checking organizations. All of the respondents performed fact-checking professionally as part of their full-time job; none of them was employed in social media. We involved both external and in-house professionals in management positions (editors-in-chief and project coordinator) and fact-checking-only positions, but also fluid and overlapping fact-checking roles with the roles of journalist, PR manager, editorial manager, and senior research fellow. Our sample represents one of the low-resource language groups, under-explored in AI-research. The sample covers a variety of sizes and types of organizations – from small NGOs to large news agencies. The majority of organizations also collaborated with social media providers (e.g., Meta). Most of them were part of the IFCN network, which prohibits any kind of political connection. Two organizations were not members of the IFCN (and did not collaborate with social media). These organizations focused on the fact-checking of political discussions in mass media (mainly TV) and partisan news. The product of these two organizations was an article summarizing the main arguments in a misleading topic of interest (not a structured fact-check of a claim).
For this work, we define Central Europe as the countries of the Visegrád Group (Tiersky et al., 2004). While Central Europe geographically belongs to the European region, its historical and linguistic context differs from Western Europe (Eastern Block legacy, Slavic languages). This has many consequences on its current social and cultural setting and related issues (e.g., the level of Russian influence). Contextual factors affect the choices of action and the use of sources and channels in the online environment (Agarwal, 2018). We can thus assume that some needs of Central European fact-checkers can differ from the fact-checkers elsewhere in Europe/the world. In particular, we can expect differences when compared with high-resource language areas.
@@
All fact-checkers participating in the survey had to be members of recognized professional groups. Therefore, organizations from CEDMO (Central European Digital Media Observatory) hub [^20] https://cedmohub.eu/ were selected for the in-depth interviews. The survey was directed at EDMO (European Digital Media Observatory) hubs [^21] https://edmo.eu/edmo-hubs/ members and to the International Fact-Checking Network (IFCN’s) [^22] https://ifcncodeofprinciples.poynter.org/signatories signatories operating in Europe. The survey respondents were recruited through contacts provided by the EDMO or LinkedIn Premium account. Both external fact-checkers and in-house professionals were involved, but none of them was employed in social media. The position names of the respondents mentioned in their LinkedIn profiles were fact-checker, verification officer, journalist, redactor, or analyst in the fact-checking parts of the organizations.
@@ 3.3. Method 1: Interviews with fact-checkers
The semi-structured interviews were conducted with one organization at a time, and each required between one and one and a half hours. The interviews were conducted from December 2021 to February 2022 using a video conference system using the English, Slovak, or Czech language.
When designing the interview, we faced many ambiguities in the fact-checking process. The number of stages, as well as terms used to name them, varies across the literature and tends to depend on the view that the research is dedicated to (social sciences versus AI) as shown in Fig. 3. Existing work in AI typically recognized four core fact-checking stages, while the first study of today’s practices of fact-checkers (noa, 2020) mentioned three different stages. Therefore, we were obliged to redefine the stages found in the literature when designing the interview questions.
@@ 3.4. Method 2: Validation survey
The interviews yielded many problems as well as some implicit inputs on what is desired and needed for fact-checkers to be more efficient in tackling disinformation. These inputs were discussed, using brainstorming methods, and problems that were most cognitively demanding, tedious, time-demanding, and complex, as well as brought inconveniences and uncertainty with them, were selected. Finally, 13 problems were chosen for the validation survey based on the user activities and current technical feasibility, resource feasibility, and sustainability of potential AI solutions (tools) (see the validation survey questions in Annex B).
The survey also served to validate the individual responses; therefore, for each activity, the respondents were queried on:
  • how often fact-checkers need to be involved in such activity (on a six point Likert scale ranging from “more times a day”=6 to “never”=1) and
  • how difficult the activity is (on a five point Likert scale ranging from “No problem, I like to do it”=1 to “I perceive it as a big problem. I really need help with this”=5).
@@
where $C(a_{i})$ is an importance coefficient that quantifies the need for automation within activity $i$, $med(f_{i})$ is a median score of the frequency of the fact-checkers’ involvement in activity $i$ and $med(l_{i})$ is a median score of the level of difficulty of activity $i$. As such, the most severe problem would get a coefficient of 30, and the least important problem a coefficient of 1.
@@ 3.5. Results presentation
The results of the interviews, i.e., the details of the fact-checking process, based on the fact-checkers’ answers in the interviews, are summarized and visualized with conceptual models using CmapTools [^23] https://cmap.ihmc.us/. They are also compared with the fact-checking process of the organizations in (Micallef et al., 2022), fact-checkers’ needs in (noa, 2020; Juneja and Mitra, 2022), and utilized tools in (Dierickx, 2022). Differences are outlined in different fonts according to the legend under each figure.
The identified problems may vary across the different tasks that fact-checkers are involved in. This may be considered a limitation of our study and prompts future research focused on complex process analyses of individual fact-checking organizations.
@@ 4. Results and findings: Interviews with fact-checkers
In general, regarding RQ1, the results of the interviews indicate that the fact-checking process in Central Europe is similar to that in the rest of the world. However, since the fact-checking process differs in the various research studies (see Fig. 3) and some of the stages were missing according the results of our content analysis, we renamed and extended its stages from the previous works (see Fig. 3). Stage 2 (Selection of potential false claims/narratives) and Stage 4 (Verification of content credibility and veracity, in computer science labeled as evidence retrieval and claim verification) were inspired by (Barrón-Cedeño et al., 2020; Nakov et al., 2021a), but based on the interviews with fact-checkers, the names of the stages were renamed (extended) following these findings: 1) fact-checkers do not select only claims, but also narratives, 2) fact-checkers verify the sources of claims besides individual claims (collectively referred to as content).
We expanded the fact-checking process as it has been understood in previous AI research in: Stage 1 (Monitoring the online space), Stage 3 (Communication and avoiding duplication), and Stage 5 (Dissemination of fact-checks). We also extended the dissemination part of the process: no longer it means only the publication and marketing of fact-checks, but includes a novel subcategory – the identification of the other versions (i.e., other appearances in the online content) of the same misinformation. We use the term “misinformation”, as not every respondent was fact-checking at the level of claims: some of them were interested in repetitive topics (narratives), whereas their final product was a fact-checking article (not a structured fact-check of a claim). This is in accordance with two identified types of fact-checking (Juneja and Mitra, 2022): 1) short-term claims centric and 2) long-term advocacy centric.
As we found out, some identified problems were specific to the investigated region and connected mainly to the capacity and financial reasons (e.g., weaker collaboration with other fact-checkers or dissemination in an attractive form) and limited language capabilities of the tools (e.g., for monitoring potential misinformation or for finding the previously published fact-checks). The virtue of Central European fact-checkers was ingenuity when forced to use the freely available tools. Its manifestation could be seen in the use of personalization on social networks using fake profiles to monitor misinformation, but also in the creative use of CrowdTangle to monitor notorious liars. Below we present a detailed description of the process and problems of our respondents in comparison with the fact-checkers around the world.
@@ 4.1. Monitoring the online space
“Monitoring of online space is done in two ways: manually or automated. When authors browse media manually (the classic media such as the New York Times and the Washington Post), they search for topics that are interesting for our audience. Some of us also do automated monitoring in CrowdTangle… I do that every morning by coffee, where I check notorious sources – some lists have about 800 [social media] pages.“
@@
Monitoring is a part of the process that fact-checkers are the most open to autonomating (Dierickx, 2022). Still, there is just a limited pool of tools that are a bit helpful for Central European as well as other non-English-speaking fact-checkers on social media.
For Meta’s platforms (Facebook, Instagram), fact-checkers originally used to rely on CrowdTangle. As this tool was originally developed for marketing purposes, it has put overemphasis on virality, which was not always useful for fact-checkers. This was mentioned by our respondents as well as other fact-checkers around the world (noa, 2020; Dierickx, 2022; Micallef et al., 2022). Despite this, fact-checkers agreed on the usefulness of this tool for creating and monitoring lists of notorious liars. It was also useful for checking the popularity of misinformation, for example, in cases when it was reported by regular users. Despite a clear value provided to fact-checkers, Meta abandoned its active development at first, and in August 2024 it was deprecated completely. From that time, fact-checkers are provided with a replacement tool built on the top of new Meta Content Library (MCL), which, however, does not provide the same set of features as the original CrowdTangle tool and which are recognized as missing ones by fact-checkers.
In the case of the original Twitter platform, fact-checkers relied on a simple but efficient TweetDeck tool. Its current version, branded as X Pro, became in July 2023 a paid proprietary tool that is available only to verified and premium users of the X platform, which reduced its availability and usefulness for fact-checkers.
In Norway, fact-checkers rather rely on their Storyboard tool to find problematic content (Dierickx, 2022), in other research  (Juneja and Mitra, 2022) BuzzSumo, Social searcher or Influencer are mentioned.
@@ 4.2. Selection of potential false claims/narratives
Some Central European fact-checkers also mentioned timeliness, factuality, representativeness, controversiality of potential misinformation, or sympathy of the topic to the authors as factors for selection. All these factors are consistent with previous research (Micallef et al., 2022), as well as putting more emphasis on societal impact than virality, including issues that affect minorities or disadvantaged groups. In comparison to our respondents, the volume of reader requests, verifiability, and the importance of providing a different perspective for their readers were mentioned as important factors for prioritization in previous research (Micallef et al., 2022). Some respondents in this research preferred to fact-check influencers, while others were aware of the high probability of a backlash from their followers and preferred not to check. Note that fact-checkers all over the world face legal actions. Some even reported facing personal danger, even physical attacks (noa, 2020; Juneja and Mitra, 2022). The latter has not been the case of our respondents, but it underlines a need for a tool that alerts fact-checkers before something goes viral.
Checking whether a claim is verifiable is one of the tasks that the fact-checkers are open to automating the most (Dierickx, 2022). In some countries, it is performed by a simple Google search or by using freely available tools  (Dierickx, 2022) like ClaimBuster (Hassan et al., 2017), BSDetector (Micallef et al., 2022) or Full Fact AI. There are also open-source tools developed and used by fact-checking organizations in Latin America that: 1) differentiate between the opinion and factual statements (Beers, 2019); and 2) optimize the process of collecting and responding to WhatsApp reader requests (Bhuiyan et al., 2020). None of these tools was mentioned or utilized by our interviewees. Study (Dierickx, 2022) mentions five reasons for not adopting AI tools by fact-checkers: 1) lack of knowledge about the existence of such tools; 2) lack of time to learn the tool; 3) lack of resources to pay for the tool; 4) lack of knowledge of how the AI tool works; and 5) difficulty in integrating the tool into the organization’s workflow. Keeping fact-checkers in the design process of the AI tools and appropriate training may help overcome these barriers (Dierickx, 2022).